# Do MLLMs Capture How Interfaces Guide User Behavior? A Benchmark for Multimodal UI/UX Design Understanding

### v3

@@ E.1 Setup
  • Claude 3.5 Sonnet (Anthropic 2024):
@@
E.2 UI/UX Design Selection
Prompts
Figures A through G show the prompts used, including those designed for diverse reasoning strategies.
Token Counts
Token counts across reasoning strategies on proprietary models are presented in Table C.
Standard Deviation
Table D presents the standard deviation across three experiment runs for each model and reasoning method. For metric abbreviation, FS and SS denote the standard deviation when the first or second UI image input was the more effective one, while AS and CS refer to the standard deviation of total correct and both correct cases, respectively.
Additional Qualitative Results
Figure N and O present the additional qualitative examples across different models. Figure P shows an example comparing various reasoning strategies executed with GPT-4o.
E.3 UI/UX Design Rationale Alignment
Prompts
Figure H presents the prompt used for models’ rationale generation, and Figure I presents the prompt used for its GPT-4o-based evaluation.
Results on Text Similarity Metrics
Alongside rationale-wise and instance-wise accuracies from GPT-4o-based evaluation, we report BLEU, ROUGE, METEOR, and BERTScore as well. (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2020) in Table E.
Additional Qualitative Results
Figure Q presents an additional qualitative example across different models.
Appendix F Human Study Details
Figure M displays the interface used for human evaluation, where participants select which synthetic UI better adheres to the given instructions, including the specified behavioral goal. Table F presents the results of the human study, showing the distribution of participant responses across 10 A/B questions. For each question, the option chosen by the majority is treated as a proxy for human preference and is highlighted in bold. Results of standard deviation across three experiment runs are shown in Table G. Figure R provides an additional qualitative example of how models performed in identifying the preferred UI variant.
Appendix G Icon Attribution
@@
Perception Law of Common Region (Koffka 1922) Elements sharing a boundary or background are perceived as a unified group. This can be a visible border, a colored box, or a distinct region in the layout.
@@
Law of Proximity (Koffka 1922) Items positioned close together are seen as belonging to the same group. This mental shortcut helps users quickly identify relationships between adjacent elements.
@@
Law of Prägnanz (Koffka 1922) Humans naturally perceive and interpret complex images in their simplest forms. It reduces cognitive effort by helping us group and understand elements more quickly.
@@
Law of Similarity (Koffka 1922) People group together elements that look alike, assuming they serve a similar purpose. Consistency in color, shape, or size leads the user to see these items as related.
@@
Memory Von Restorff Effect (Von Restorff 1933) People remember a distinct element more easily than others that blend in. This is also known as the Isolation Effect, emphasizing our attention on anything unusual.
@@
Serial Position Effect (Ebbinghaus, Ruger, and Bussenius 1913) Users tend to remember the first and last items in a series more than those in the middle. This influences how they recall sequential information or actions.
@@
Miller’s Law (Miller 1956) On average, people can hold only about seven items in their working memory at once. Going beyond this limit causes cognitive overload.
@@
Jakob’s Law (Nielsen 2000) Users expect new interfaces to operate similarly to those they’ve used before. Leveraging established patterns lowers cognitive load and accelerates adoption.
@@
Action Hick’s Law (Hick 1952) Decision time rises with the number and complexity of choices. Too many options can overwhelm users and slow their actions.
@@
Pareto Principle (Pareto 1919) Roughly 80% of outcomes stem from 20% of causes. Focusing on the most impactful features or tasks delivers the greatest improvement.
@@
Fitts’s Law (Fitts and Peterson 1964) The time required to reach a target depends on its size and distance. Larger, closer elements are easier to access quickly and accurately.
@@
Goal-Gradient Effect (Hull 1932) Motivation increases as people get closer to finishing a task. Providing clear progress markers encourages users to continue and complete it faster.
@@
Table A: Overview of UX laws used for rationale annotation, and their categorization by cognitive dimension (Perception, Memory, Action).
Type Definition
Content & Copy All textual elements that convey meaning or guide the user, including headlines, body text, labels, error messages, promotional language, and social proof content.
Visual & Style Visual design elements such as color schemes, typography, spacing, icon sizes, and overall aesthetic consistency that shape the look and feel of the interface.
Controls & Inputs Interactive UI components that allow users to perform actions, including buttons, toggles, dropdowns, sliders, search fields, and other input elements.
Layout & Structure The spatial organization of UI components, covering alignment, grouping, column or grid usage, visual hierarchy, and structural consistency across screens.
Navigation & Flow Interface elements that support user movement and orientation within the product, such as menus, sidebars, breadcrumbs, tabs, and filter panels.
Media & Visual Assets Non-textual content like images, icons, thumbnails, banners, and videos used to support communication, branding, or user engagement.
Pricing & Promotion How pricing details, discounts, promotional messages, limited-time offers, or benefits like free trials are visually and textually communicated to the user.
Trust & Social Proof UI elements that aim to build user trust and credibility, including customer reviews, ratings, trust badges, security seals, guarantees, and testimonials.
Forms & Form UX Design and usability aspects of data entry forms, such as number and arrangement of fields, required indicators, input validation, and overall ease of completion.
Status & Feedback Visual and textual responses to user actions or system states, including loading indicators, progress bars, tooltips, error messages, confirmations, and alerts.
Table B: Definitions of UI element change types.
@@
Table C: Overview of Token Counts ($\mu\pm\sigma$) in the UI/UX design selection task on WiserUI-Bench.
Model Method FS SS AS CS
@@
Table D: Results on standard deviation across three experiment runs in the UI/UX design selection task for each model and reasoning strategy on WiserUI-Bench.
@@
Table E: Results of text similarity metrics per expert rationale for the UI/UX design rationale alignment task on WiserUI-Bench.
Choice Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q10
A 42 9 23 14 25 40 14 27 35 18
B 13 46 32 41 30 15 41 28 20 37
Majority Choice A B B B B A B B A B
Table F: Distribution of responses from human participants indicating their preferred UI variant across 10 A/B test questions.
Model FS SS AS CS
GPT-4o 0.577 0.000 0.577 0.577
Claude 3.5 Sonnet 0.577 0.577 0.577 0.577
Qwen2.5-VL-32B 0.577 0.577 1.000 0.000
InternVL-2.5-38B 0.577 0.577 0.577 1.000
Table G: Results on standard deviation across three experiment runs on identifying the behaviorally preferred synthetic UI variant in human study.
@@
Figure A: Prompt used for the UI/UX design selection task with zero-shot strategy.
@@
Figure B: Prompt used for the UI/UX design selection task with CoCoT strategy.
@@
Figure C: Prompt used for the UI/UX design selection task with Self-Refine strategy on the reviewing step. Prompt inputs are in bold.
@@
Figure D: Prompt used for the UI/UX design selection task with Self-Refine strategy on the improving step. Prompt inputs are in bold.
@@
Figure E: Prompt used for the UI/UX design selection task with DDCoT strategy.
@@
Figure F: Prompt used for the UI/UX design selection task with MAD strategy on the debate step arguing the first/second version better drives user actions. Prompt inputs are in bold.
@@
Figure G: Prompt used for the UI/UX design selection task with MAD strategy on the moderation step. Prompt inputs are in bold.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. If the SECOND version has proven to be more effective UI/UX design that leads to better user experience and conversion, provide all possible and appropriate for its superiority.
Figure H: Prompt used for the UI/UX design rationale alignment task.
Given two statements are reasons why one UI is better than another in terms of guiding user behavior effectively. Statement 1: {model_reason}
@@
Figure I: Prompt used for GPT-4o-based evaluation on the UI/UX design rationale alignment task. Prompt inputs are in bold.
Figure J: Example of visual indicator removal in WiserUI-Bench. Left: Original image with markers added by data source website. Right: After inpainting, with markers removed while preserving UI elements.
Figure K: Detailed statistics of WiserUI-Bench on (a) page type, (b) industry domain, (c) device type, (d) UI element change type
Figure L: A screenshot of the rationale annotation interface for UI/UX experts based on original data sources and guidelines. While writing the key rationales, they also selected the corresponding UX law.
Figure M: A screenshot of the human study interface. Each participant selects the UI variant they believe better supports the given instruction, including the specified behavioral goal.
Figure N: A qualitative example from the UI/UX design selection task on WiserUI-Bench: Both Claude 3.5 Sonnet and InternVL-2.5-38B selected the A/B test winner.
Figure O: A qualitative example from the UI/UX design selection task on WiserUI-Bench: GPT-4o selected the A/B test winner, whereas LLaVA-NeXT-7B did not.
Figure P: A qualitative example from the UI/UX design selection task on WiserUI-Bench, showcasing the effect of diverse reasoning strategies executed on GPT-4o.
Figure Q: A qualitative example from the UI/UX design rationale alignment task on WiserUI-Bench: InternVL-2.5-8B captured only the memory rationale, while Qwen2.5-VL-32B identified two key rationales. The highlights in the models’ responses indicate the parts that align with the expert-curated key rationales.
Figure R: A qualitative example from the UI/UX design selection on human study: InternVL2.5-38B failed to match the majority choice, whereas Claude 3.5 Sonnet succeeded. For fair evaluation, we inserted the same images into the placeholder area of each generated UI, as shown above.

### v4

@@ B.1.1 Model Configuration
  • Claude 3.5 Sonnet Anthropic (2024):
@@ B.1.2 Resources
B.2 UI/UX Design Selection
B.2.1 Prompts
Figures 8 through 14 show the prompts used, including those designed for diverse reasoning strategies.
B.2.2 Token Counts
Token counts across reasoning strategies on proprietary models are presented in Table 6.
B.2.3 Standard Deviation
Table 7 presents the standard deviation across three experiment runs for each model and reasoning method. For metric abbreviation, FS and SS denote the standard deviation when the first or second UI image input was the more effective one, while AS and CS refer to the standard deviation of total correct and both correct cases, respectively.
B.2.4 Additional Case Study
Figure 21 and 22 present the additional case studies across different models. Figure 23 shows an example comparing various reasoning strategies executed with GPT-4o.
B.3 UI/UX Design Interpretation
B.3.1 Prompts
Figure 15 presents the prompt used for models’ interpretation generation, and Figure 16 presents the prompt used for its GPT-4o-based evaluation.
B.3.2 Results on Text Similarity Metrics
Alongside Interpretation Recall and Instance-level Recall from GPT-4o-based evaluation, we report BLEU, ROUGE, METEOR, and BERTScore as well. Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005); Zhang et al. (2020) in Table 8.
B.3.3 Additional Case Study
Figure 24 presents an additional case study across different models.
Appendix C Icon Attribution
@@
Perception Law of Common Region Koffka (1922) Elements sharing a boundary or background are perceived as a unified group. This can be a visible border, a colored box, or a distinct region in the layout.
@@
Law of Proximity (Koffka, 1922) Items positioned close together are seen as belonging to the same group. This mental shortcut helps users quickly identify relationships between adjacent elements.
@@
Law of Prägnanz (Koffka, 1922) Humans naturally perceive and interpret complex images in their simplest forms. It reduces cognitive effort by helping us group and understand elements more quickly.
@@
Law of Similarity (Koffka, 1922) People group together elements that look alike, assuming they serve a similar purpose. Consistency in color, shape, or size leads the user to see these items as related.
@@
Memory Von Restorff Effect (Von Restorff, 1933) People remember a distinct element more easily than others that blend in. This is also known as the Isolation Effect, emphasizing our attention on anything unusual.
@@
Serial Position Effect (Ebbinghaus et al., 1913) Users tend to remember the first and last items in a series more than those in the middle. This influences how they recall sequential information or actions.
@@
Miller’s Law (Miller, 1956) On average, people can hold only about seven items in their working memory at once. Going beyond this limit causes cognitive overload.
@@
Jakob’s Law (Nielsen, 2000) Users expect new interfaces to operate similarly to those they’ve used before. Leveraging established patterns lowers cognitive load and accelerates adoption.
@@
Action Hick’s Law (Hick, 1952) Decision time rises with the number and complexity of choices. Too many options can overwhelm users and slow their actions.
@@
Pareto Principle (Pareto, 1919) Roughly 80% of outcomes stem from 20% of causes. Focusing on the most impactful features or tasks delivers the greatest improvement.
@@
Fitts’s Law (Fitts and Peterson, 1964) The time required to reach a target depends on its size and distance. Larger, closer elements are easier to access quickly and accurately.
@@
Goal-Gradient Effect (Hull, 1932) Motivation increases as people get closer to finishing a task. Providing clear progress markers encourages users to continue and complete it faster.
@@
Table 5: Overview of UX laws used for interpretation annotation, and their categorization by cognitive dimension (Perception, Memory, Action).
@@
Table 6: Overview of Token Counts ($\mu\pm\sigma$) in the UI/UX design selection task on WiserUI-Bench.
Model Method FS $S$ AS CS
@@
Table 7: Results on standard deviation across three experiment runs in the UI/UX design selection task for each model and reasoning strategy on WiserUI-Bench.
@@
Table 8: Results of text similarity metrics per expert interpretation for the UI/UX design interpretation task on WiserUI-Bench.
@@
Figure 8: Prompt used for the UI/UX design selection task with zero-shot strategy.
@@
Figure 9: Prompt used for the UI/UX design selection task with CoCoT strategy.
@@
Figure 10: Prompt used for the UI/UX design selection task with Self-Refine strategy on the reviewing step. Prompt inputs are in bold.
@@
Figure 11: Prompt used for the UI/UX design selection task with Self-Refine strategy on the improving step. Prompt inputs are in bold.
@@
Figure 12: Prompt used for the UI/UX design selection task with DDCoT strategy.
@@
Figure 13: Prompt used for the UI/UX design selection task with MAD strategy on the debate step arguing the first/second version better drives user actions. Prompt inputs are in bold.
@@
Figure 14: Prompt used for the UI/UX design selection task with MAD strategy on the moderation step. Prompt inputs are in bold.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. If the SECOND version has proven to be more effective UI/UX design that leads to better user experience and conversion, provide all possible and appropriate interpretations for its superiority.
Figure 15: Prompt used for the UI/UX design interpretation task.
Given two statements are interpretations why one UI is better than another in terms of guiding user behavior effectively. Statement 1: {model_reason}
@@
Figure 16: Prompt used for GPT-4o-based evaluation on the UI/UX design interpretation task. Prompt inputs are in bold.
Figure 17: Example of visual indicator removal in WiserUI-Bench. Left: Original image with markers added by data source website. Right: After inpainting, with markers removed while preserving UI elements.
Figure 18: Detailed statistics of WiserUI-Bench on (a) page type, (b) industry domain, (c) device type
Figure 19: Detailed statistics of WiserUI-Bench on UI change attribute type
Figure 20: A screenshot of the interpretation annotation interface for UI/UX experts based on original data sources and guidelines. After writing the key interpretations, they also selected the most relevant UX law.
Figure 21: A qualitative example from the UI/UX design selection task on WiserUI-Bench: Both Claude 3.5 Sonnet and InternVL-2.5-38B selected the A/B test winner.
Figure 22: A qualitative example from the UI/UX design selection task on WiserUI-Bench: GPT-4o selected the A/B test winner, whereas LLaVA-NeXT-7B did not.
Figure 23: A qualitative example from the UI/UX design selection task on WiserUI-Bench, showcasing the effect of diverse reasoning strategies executed on GPT-4o.
Figure 24: A qualitative example from the UI/UX design interpretation task on WiserUI-Bench: InternVL-2.5-8B captured only the memory part, while Qwen2.5-VL-32B identified two key interpretations. The highlights in the models’ responses indicate the parts that align with the expert-curated key interpretations.